# LTM: Scalable and Black-box Similarity-based Test Suite Minimization based on Language Models

### v3

@@ Abstract
Test suites tend to grow when software evolves, making it often infeasible to execute all test cases with the allocated testing budgets, especially for large software systems. Test suite minimization (TSM) is employed to improve the efficiency of software testing by removing redundant test cases, thus reducing testing time and resources, while maintaining the fault detection capability of the test suite. Most existing TSM approaches rely on code coverage (white-box) or model-based features, which are not always available to test engineers. Recent TSM approaches that rely only on test code (black-box) have been proposed, such as ATM and FAST-R. To address the scalability, we propose LTM (Language model-based Test suite Minimization), a novel, scalable, and black-box similarity-based TSM approach based on large language models (LLMs), which is the first application of LLMs in the context of TSM. To support similarity measurement for test code embeddings, we investigate five pre-trained language models: CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama, on which we compute two similarity measures: Cosine Similarity and Euclidean Distance. Our goal is to find similarity measures that are not only computationally more efficient but can also better guide a Genetic Algorithm (GA) to search for optimal minimized test suites, thus reducing the overall search time. Experimental results show that the best configuration of LTM (UniXcoder/Cosine) outperforms ATM in three aspects: (a) achieving a slightly greater saving rate of testing time ($41.72\%$ versus $40.29\%$, on average); (b) attaining a significantly higher fault detection rate ($0.84$ versus $0.81$, on average); and, most importantly, (c) minimizing test suites nearly five times faster on average, with higher gains for larger test suites and systems, thus achieving much higher scalability.
@@ 1 Introduction
Software testing is an essential activity that ensures high-quality software systems by detecting faults in software releases [1]. When test suites are large, especially for large code bases, the scalability of running all test cases in a test suite becomes quickly a critical issue, as time and resources are limited for testing in industrial contexts [2]. The problem is particularly acute as test suites tend to increase in size over time with software evolution, making it impossible to run all test cases for every code change with allocated testing budgets [1]. To address this issue, test suite minimization (TSM) is widely employed to make the testing process more efficient while maintaining its fault detection capability [1]. In contrast to test case selection [3], which selects a subset of test cases that are relevant to code changes, and test case prioritization [3], which ranks test cases based on certain characteristics, such as fault detection capability, the objective of test suite minimization is to eliminate redundant or similar test cases that are unlikely to detect different faults [3, 1].
@@
Though there exist various TSM approaches, most of them rely on either code coverage information (white-box) [3], which requires access to software production code, or model-based features, such as transition coverage for UML state machine-based testing [2]. This information is not always accessible by test engineers and can be challenging to collect [4, 5]. Further, collecting code coverage information is expensive [6, 7] and can lead to scalability issues for large software systems. Cruciani et al. [6] and Pan et al. [8] addressed this issue by proposing TSM approaches that rely only on the source code of test cases (black-box). FAST-R, which is based on clustering algorithms, has been shown to be much more efficient than white-box approaches, but achieved relatively low fault detection rates. ATM, which is based on test case similarity and evolutionary search, achieved higher fault detection rates within practically acceptable, yet longer time. Since test suite minimization is performed on major software releases [9, 8], as opposed to test case selection and prioritization which are performed for every code change, ATM provides a better trade-off compared to FAST-R in many practical situations. However, ATM still presents scalability issues for very large software systems, since its total minimization time (the sum of preparation time and search time) increases rapidly with test suite size, represented by the number of test cases per version. Our analysis of the total minimization time of ATM reveals that similarity measures play a major role in limiting its scalability, due to the fact that (1) computing tree-based similarity is expensive, taking up to $41.20\%$ of the total minimization time, and (2) similarity measures impact the search convergence and speed. This motivated us to investigate similarity measures that are both more efficient to calculate and more informative to guide the search.
@@
Language models pre-trained on programming languages convert test code into vector-based embeddings. This enables vector-based similarity measurement, which facilitates implementation optimizations and results in much higher computational efficiency than tree-based similarity measurement that relies on traversing AST trees. Moreover, these language models are pre-trained on large source code corpora with various code understanding and generation tasks, thus generating embeddings that capture informative syntactic and contextual details from test code. This suggests that such embeddings with vector-based similarity might be potentially more informative than tree-based similarity in guiding evolutionary search. Therefore, in this paper, we propose LTM (Language model-based Test suite Minimization), a scalable, black-box similarity-based TSM approach that is based on pre-trained Large Language Models (LLMs) and vector-based similarity measures, making it the first application of LLMs in the context of TSM, to the best of our knowledge. LTM uses the source code of test cases (Java test methods), without requiring any preprocessing, as input to five alternative pre-trained language models, namely CodeBERT [10], GraphCodeBERT [11],UniXcoder [12], StarEncoder [13], and CodeLlama [14], in order to extract test method embeddings. Considering that test suite diversity was reported to have a positive correlation with fault detection [2, 15, 16], LTM employs two similarity measures: Cosine similarity and Euclidean distance, to calculate the similarity between the extracted embeddings. Using the calculated similarity values, LTM employs a Genetic Algorithm (GA) to minimize test suites, which searches for the most optimal subset of a test suite, for a given testing budget. We evaluated LTM with the same dataset used to evaluate ATM (Defects4J), followed a similar experimental design, and used the same evaluation metrics of effectiveness and efficiency: Fault Detection Rate (FDR) and Minimization Time (MT), respectively. In addition, considering the potential variation in test case execution times, assessing minimized test suites based on their number of test cases only might not always be accurate [1]. Therefore, we extended our evaluation of minimized test suites using an additional metric: Time Saving Rate (TSR). Moreover, we optimized GA by utilizing a more efficient data structure to accelerate fitness calculation and enhance memory usage, which led to a $273$-fold reduction in minimization time, without requiring any additional computation resources. Then, we identified the best configuration of LTM by considering both effectiveness and efficiency, among all its alternatives, and compared it to the two best ATM configurations. Finally, we expanded our experiments by running LTM on a much larger project that ATM could not handle, to further assess the scalability of our approach.
@@
  • RQ1: How does LTM perform for test suite minimization under different configurations? LTM achieves high FDR results (an overall average FDR of $0.79$ across configurations) for a $50\%$ minimization budget (i.e., the percentage of test cases retained in the minimized test suite). The best configuration of LTM is UniXcoder using Cosine similarity when considering both effectiveness ($0.84$ FDR on average) and efficiency ($0.82$ min on average), which also achieves a greater time saving rate (an average TSR of $41.72\%$). For the large project, $Closure$, UniXcoder using Cosine Similarity takes only $17.90$ min in terms of MT and achieves an FDR of $0.79$, while saving $52.55\%$ of testing time.
@@
  • We optimize the search process of LTM by utilizing a more efficient data structure for fitness calculation, which in turn reduced the search time by $273$ folds.
@@ 2 Related Work
Test suite minimization (TSM) aims at improving the efficiency of software testing by removing redundant test cases, thus reducing testing time and resources while maintaining the effectiveness of the test suite (i.e., fault detection capability) [1]. There are various approaches that have been proposed to address it, including (a) greedy heuristics-based approaches [17, 9], which select test cases iteratively based on code coverage information, (b) clustering-based approaches [18, 19], which group test cases based on the similarity of their coverage information, and (c) search-based approaches [20, 2], which employ evolutionary search to find an optimal subset of the test suite using model-based features. Most of these approaches utilize code coverage information (white-box) [3] (i.e., requires access to the system production code) or model-based features, which are not always available to test engineers [8, 5]. Moreover, collecting code coverage information can result in up to $30\%$ time overhead [6, 7], making it not scalable for large test suites and systems.
Viggiato et al. [21] investigate the similarity of test cases that are written in natural language. Though their objective is to identify redundancy among test cases, as opposed to TSM within a budget, their similarity measures can priori be further used for this purpose. They convert test cases into vector-based representations using five text embedding techniques, i.e., Word2Vec [22], BERT [23], Sentence-BERT [24], Universal Sentence Encoder [25] and TF-IDF (Term Frequency–Inverse Document Frequency) [26]. They then identify similar test cases utilizing clustering algorithm to group similar test steps based on two different similarity metrics (Word Mover’s Distance (WMD) [27] and Cosine Similarity). The results show that their approach achieves an F-Score of $83.47\%$ for identifying similar test cases. In summary, different from our work, their focus is on identifying redundant test cases written in natural language, as opposed to code, and they do not aim to optimize a test suite within a budget.
@@
However, ATM suffers from limitations regarding its scalability for very large projects in the dataset, such as $Time$, which consists of $30k$ lines of code in its most recent version with large test suites (nearly $4k$ test cases per version), as minimization took more than $10$ hours per version as compared to FAST-R ($2.17$ seconds) and random minimization ($0.006$ seconds). The similarity measures used by ATM had a direct impact on its scalability. First, the similarity calculation time took up to $41.20\%$ (using tree edit distance) of the overall minimization time of ATM, which is due to the fact that converting test code into ASTs and calculating tree-based similarity based on ASTs are both time- and resource-consuming. Second, the search time, which was influenced by the employed similarity measures and the number of test cases per version, increased rapidly with the test suite size. Interestingly, though tree edit distance took much longer to calculate than other ATM similarity measures, it enabled GA search to converge faster towards the termination criterion (a fitness improvement of less than $0.0025$ across generations). Although GA with combined similarity was identified as the best ATM configuration based on the average minimization time across projects ($1.2~{}hours$), it is important to note that it was $2.5~{}hours$ slower than GA with tree edit distance for the largest test suite with nearly $4k$ test cases. Therefore, when dealing with large test suites, adopting a more informative similarity measure that facilitates faster search convergence can result in a notable reduction in minimization time, regardless of the computational cost associated with the similarity calculation.
In contrast to the above TSM approaches, LTM relies on test code without requiring any preprocessing, and employs large pre-trained language models (i.e., CodeBERT, GraphCodeBERT, and UniXcoder) and commonly used vector-based similarity measures (i.e., Cosine similarity and Euclidean distance) [33] to generate more informative similarity values between test case pairs. The goal is to both calculate similarity more efficiently and better guide the search algorithm (GA) to converge faster to a better fault detection capability.
@@ 3.1 Language Models for Test Method Representation
  • Unlike the above models that are pre-trained for code understanding tasks, CodeLlama is specialized for code generation and infilling tasks (i.e., filling a missing part of a code snippet). It has $7Billion$ parameters, which is significantly higher than the parameter size ($125Million$) of the above models, and was pre-trained on a larger dataset, making it potentially able to distinguish test code more precisely.
@@ 3.1.2 Generating Test Method Embeddings
  • GraphCodeBERT extend such architecture using a graph-guided masked attention function to help the model encode graph-based data flow information.
@@
  • Unlike the above models, CodeLlama employs code infilling task, which allows the model to generate the missing part of a code snippet while comprehending the entire context. This enables the model to understand the logical structure of the test code and the relationships between test code segments.
@@ 4.1 Research Questions
  • RQ1.1: How effectively can LTM minimize test suites? The language models used for LTM (i.e., CodeBERT, GraphCodeBERT, and UniXcoder) vary in terms of architectures and training data, thus generating embeddings that capture different information about test case code. The distance functions we used to measure the similarity or distance between test method embeddings differently, which can lead to different comparison results. Therefore, the combination of language models and distance functions can affect the ability of LTM to remove redundant test cases, which can in turn affect its fault detection capability. In this RQ, we assess the effectiveness of LTM in terms of fault detection capability under different configurations. In addition, we calculated the execution time of minimized test suites to quantify the achieved time savings resulting from test suite minimization.
@@ 4.2.1 Baseline approach (ATM)
We compared LTM to ATM [8], a similarity-based, search-based test suite minimization approach that relies solely on the source code of test cases. ATM preprocesses test code by removing information that is irrelevant to testing logic from the test code, then converts test code into ASTs. ATM calculates four different tree-based similarity measurements between test case ASTs: top-down, bottom-up, combined and tree edit distance similarity. Search algorithms (i.e., GA and NSGA-II) were then employed based on similarity values to find an optimal subset of test cases by removing redundant test cases. Using GA with combined similarity, ATM achieved a better trade-off in terms of effectiveness and efficiency compared to a SOTA test suite minimization approach, FAST-R [6]. However, it suffered from scalability issues as, for example, it took more than $10$ hours to minimize $4k$ test cases for the largest test suite. This motivated us to compare LTM to ATM, in terms of fault detection capability and, more importantly, minimization time. We aim to assess whether LTM can be more scalable than ATM while achieving comparable or even higher fault detection capability. We compared LTM to the two best configurations of ATM, using GA with combined similarity (ATM/Combined) and tree edit distance (ATM/TreeEditDistance). The former was the best configuration considering the trade-off between effectiveness and efficiency, whereas the latter achieved higher fault detection capability but was more time-consuming than the former in terms of total minimization time. We further evaluated the achieved saving in testing time for every minimized test suite for both LTM and ATM.We relied on the publicly available replication package of ATM [^6] https://doi.org/10.5281/zenodo.7455766 provided by its authors.

### v4

@@ Abstract
Test suites tend to grow when software evolves, making it often infeasible to execute all test cases with the allocated testing budgets, especially for large software systems. Test suite minimization (TSM) is employed to improve the efficiency of software testing by removing redundant test cases, thus reducing testing time and resources while maintaining the fault detection capability of the test suite. Most existing TSM approaches rely on code coverage (white-box) or model-based features, which are not always available to test engineers. Recent TSM approaches that rely only on test code (black-box) have been proposed, such as ATM and FAST-R. The former yields higher fault detection rates (FDR) while the latter is faster. To address scalability while retaining a high FDR, we propose LTM (Language model-based Test suite Minimization), a novel, scalable, and black-box similarity-based TSM approach based on large language models (LLMs), which is the first application of LLMs in the context of TSM. To support similarity measurement using test method embeddings, we investigate five different pre-trained language models: CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama, on which we compute two similarity measures: Cosine Similarity and Euclidean Distance. Our goal is to find similarity measures that are not only computationally more efficient but can also better guide a Genetic Algorithm (GA), which is used to search for optimal minimized test suites, thus reducing the overall search time. Experimental results show that the best configuration of LTM (UniXcoder/Cosine) outperforms ATM in three aspects: (a) achieving a slightly greater saving rate of testing time ($41.72\%$ versus $41.02\%$, on average); (b) attaining a significantly higher fault detection rate ($0.84$ versus $0.81$, on average); and, most importantly, (c) minimizing test suites nearly five times faster on average, with higher gains for larger test suites and systems, thus achieving much higher scalability.
@@ 1 Introduction
Software testing is an essential activity that ensures high-quality software systems by detecting faults in software releases [1]. When test suites are large, especially for large code bases, the scalability of running all test cases in a test suite quickly becomes a critical issue, as time and resources are limited for testing in industrial contexts [2]. The problem is particularly acute as test suites tend to increase in size over time with software evolution, making it impossible to run all test cases for every code change with allocated testing budgets [1]. To address this issue, test suite minimization (TSM) is widely employed to make the testing process more efficient while maintaining its fault detection capability [1]. In contrast to test case selection [3], which selects a subset of test cases that are relevant to code changes, and test case prioritization [3], which ranks test cases based on certain characteristics, such as fault detection capability, the objective of test suite minimization is to eliminate redundant or similar test cases that are unlikely to detect different faults [3, 1].
@@
Though there exist various TSM approaches, most of them rely on either code coverage information (white-box) [3], which requires access to software production code, or model-based features, such as transition coverage for UML state machine-based testing [2]. This information is not always accessible by test engineers and can be challenging to collect [4, 5]. Further, collecting code coverage information is expensive [6, 7] and can lead to scalability issues for large software systems. Cruciani et al. [6] and Pan et al. [8] addressed this issue by proposing TSM approaches that rely only on the source code of test cases (black-box). FAST-R, which is based on clustering algorithms, has been shown to be much more efficient than white-box approaches, but achieved relatively low fault detection rates. ATM, which is based on test case similarity and evolutionary search, achieved higher fault detection rates within practically acceptable, yet longer time. Since test suite minimization is performed on major software releases [9, 8], as opposed to test case selection and prioritization which are performed for every code change, ATM offers a better trade-off compared to FAST-R in many practical situations. However, ATM still presents scalability issues for very large software systems, since its total minimization time (the sum of preparation time and search time) increases rapidly with test suite size, represented by the number of test cases per project version. Our analysis of the total minimization time of ATM reveals that similarity measures play a major role in limiting its scalability, due to the fact that (1) computing tree-based similarity is expensive, taking up to $41.20\%$ of the total minimization time, and (2) similarity measures impact the search convergence and speed. This motivated us to investigate similarity measures that are both more efficient to calculate and more informative to guide the search.
@@
Language models pre-trained on programming languages convert test code into vector-based embeddings. This enables vector-based similarity measurement, which facilitates implementation optimizations and results in much higher computational efficiency than tree-based similarity measurement that relies on traversing AST trees. Moreover, these language models are pre-trained on large source code corpora with various code understanding and generation tasks, thus generating embeddings that capture informative syntactic and contextual details from test code. This suggests that such embeddings with vector-based similarity might be potentially more informative than tree-based similarity in guiding evolutionary search. Therefore, in this paper, we propose LTM (Language model-based Test suite Minimization), a scalable, black-box similarity-based TSM approach that is based on pre-trained Large Language Models (LLMs) and vector-based similarity measures, making it the first application of LLMs in the context of TSM, to the best of our knowledge. LTM uses the source code of test cases (Java test methods), without requiring any preprocessing, as input to five alternative pre-trained language models, namely CodeBERT [10], GraphCodeBERT [11],UniXcoder [12], StarEncoder [13], and CodeLlama [14], in order to extract test method embeddings. Considering that test suite diversity was reported to have a positive correlation with fault detection [2, 15, 16], LTM employs two similarity measures: Cosine similarity and Euclidean distance, to calculate the similarity between the extracted embeddings. Using the calculated similarity values, LTM employs a Genetic Algorithm (GA) to minimize test suites, which searches for the most optimal subset of a test suite, for a given testing budget. We evaluated LTM with the same dataset used to evaluate ATM (Defects4J), followed a similar experimental design, and used the same evaluation metrics of effectiveness and efficiency: Fault Detection Rate (FDR) and Minimization Time (MT), respectively. In addition, considering the potential variation in test case execution times, assessing minimized test suites based on their number of test cases only might not always be accurate [1]. Therefore, we extended our evaluation of minimized test suites using an additional metric: Time Saving Rate (TSR). Moreover, we optimized GA by utilizing a more efficient data structure to accelerate fitness calculation and enhance memory usage, which led to a $190$-fold reduction in minimization time, without requiring any additional computation resources. Then, we identified the best configuration of LTM by considering both effectiveness and efficiency, among all its alternatives, and compared it to the two best ATM configurations. Finally, we expanded our experiments by running LTM on a much larger project that ATM could not handle, to further assess the scalability of our approach.
@@
  • RQ1: How does LTM perform for test suite minimization under different configurations? LTM achieves high FDR results (an overall average FDR of $0.79$ across configurations) for a $50\%$ minimization budget (i.e., the percentage of test cases retained in the minimized test suite). The best configuration of LTM is UniXcoder using Cosine similarity when considering both effectiveness ($0.84$ FDR on average) and efficiency ($0.82$ min on average), which also achieves a greater time saving rate (an average TSR of $41.72\%$). For the large project, $Closure$, UniXcoder using Cosine Similarity takes only $17.80$ min in terms of MT and achieves an FDR of $0.79$, while saving $52.55\%$ of testing time.
@@
  • We optimize the search process of LTM by utilizing a more efficient data structure for fitness calculation, which in turn reduced the search time by $190$ folds.
@@ 2 Related Work
Test suite minimization (TSM) aims at improving the efficiency of software testing by removing redundant test cases, thus reducing testing time and resources while maintaining the effectiveness of the test suite (i.e., fault detection capability) [1]. There are various approaches that have been proposed to address TSM, including (a) greedy heuristics-based approaches [17, 9], which select test cases iteratively based on code coverage information, (b) clustering-based approaches [18, 19], which group test cases based on the similarity of their coverage information, and (c) search-based approaches [20, 2], which employ evolutionary search to find an optimal subset of the test suite using model-based features. Most of these approaches utilize code coverage information (white-box) [3] (i.e., requires access to the system production code) or model-based features, which are not always available to test engineers [8, 5]. Moreover, collecting code coverage information can result in up to $30\%$ time overhead [6, 7], making it not scalable for large test suites and systems.
Viggiato et al. [21] investigate the similarity of test cases that are written in natural language. Though their objective is to identify redundancy among test cases, as opposed to TSM within a budget, their similarity measures can priori be further used for this purpose. They convert test cases into vector-based representations using five text embedding techniques (i.e., Word2Vec [22], BERT [23], Sentence-BERT [24], Universal Sentence Encoder [25], and TF-IDF (Term Frequency–Inverse Document Frequency) [26]). They then identify similar test cases utilizing clustering algorithms to group similar test steps based on two different similarity metrics (Word Mover’s Distance (WMD) [27] and Cosine Similarity). The results show that their approach achieves an F-Score of $83.47\%$ for identifying similar test cases. In summary, different from our work, their focus is on identifying redundant test cases written in natural language, as opposed to code, and they do not aim to optimize a test suite within a budget.
@@
However, ATM suffers from limitations regarding its scalability for very large projects in the dataset, such as $Time$, which consists of $30k$ lines of code in its most recent version with large test suites (nearly $4k$ test cases per version), as minimization took more than $10$ hours per project version as compared to FAST-R ($2.17$ seconds) and random minimization ($0.006$ seconds). The similarity measures used by ATM had a direct impact on its scalability. First, the similarity calculation time took up to $41.20\%$ (using tree edit distance) of the overall minimization time of ATM, which is due to the fact that converting test code into ASTs and calculating tree-based similarity based on ASTs are both time- and resource-consuming. Second, the search time, which was influenced by the employed similarity measures and the number of test cases per version, increased rapidly with the test suite size. Interestingly, though tree edit distance took much longer to calculate than other ATM similarity measures, it enabled GA search to converge faster towards the termination criterion (a fitness improvement of less than $0.0025$ across generations). Although GA with combined similarity was identified as the best ATM configuration based on the average minimization time across projects ($1.2~{}hours$), it is important to note that it was $2.5~{}hours$ slower than GA with tree edit distance for the largest test suite with nearly $4k$ test cases. Therefore, when dealing with large test suites, adopting a more informative similarity measure that facilitates faster search convergence can result in a notable reduction in minimization time, regardless of the computational cost associated with the similarity calculation.
In contrast to the above TSM approaches, LTM relies on test code without requiring any preprocessing, and employs large pre-trained language models (i.e., CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama) and commonly used vector-based similarity measures (i.e., Cosine similarity and Euclidean distance) [33] to generate more informative similarity values between test case pairs. The goal is to both calculate similarity more efficiently and better guide the search algorithm (GA) to converge faster to a better fault detection capability.
@@ 3.1 Language Models for Test Method Representation
  • Unlike the above models that are pre-trained for code understanding tasks, CodeLlama is specialized for code generation and infilling tasks (i.e., filling a missing part of a code snippet). It has billions of parameters, which is significantly higher than the parameter size ($125Million$) of the above models, and was pre-trained on a larger dataset, making it potentially able to distinguish test code more precisely.
@@ 3.1.2 Generating Test Method Embeddings
  • GraphCodeBERT extends such architecture using a graph-guided masked attention function to help the model encode graph-based data flow information.
@@
  • Unlike the above models, CodeLlama employs the code infilling task, which allows the model to generate the missing part of a code snippet while comprehending the entire context. This enables the model to understand the logical structure of the test code and the relationships between test code segments.
@@ 4.1 Research Questions
  • RQ1.1: How effectively can LTM minimize test suites? The language models used for LTM (i.e., CodeBERT, GraphCodeBERT, UniXcoder, StarEncoder, and CodeLlama-7b) vary in terms of architectures and training data, thus generating embeddings that capture different information about test case code. The distance functions we used to measure the similarity or distance between test method embeddings differently, which can lead to different comparison results. Therefore, the combination of language models and distance functions can affect the ability of LTM to remove redundant test cases, which can in turn affect its fault detection capability. In this RQ, we assess the effectiveness of LTM in terms of fault detection capability under different configurations. In addition, we calculated the execution time of minimized test suites to quantify the achieved time savings resulting from test suite minimization.
@@ 4.2.1 Baseline approach (ATM)
We compared LTM to ATM [8], a similarity-based, search-based test suite minimization approach that relies solely on the source code of test cases. ATM preprocesses test code by removing information that is irrelevant to testing logic from the test code, then converts test code into ASTs. ATM calculates four different tree-based similarity measurements between test case ASTs: top-down, bottom-up, combined and tree edit distance similarity. Note that ATM does not use the full code (i.e., the source code of test cases without preprocessing) as it would make the already-long preparation phase of ATM much longer and impractical in realistic settings. This is due to the fact that (1) the XML files used by ATM to store Abstract Syntax Trees (ASTs) of test cases would require substantial storage space and (2) it would take longer time to traverse the ASTs of test cases when calculating similarity, thus yielding much longer preparation time for transforming the test case code into ASTs and calculating similarity values. As a result, ATM’s preparation time can take up to 41.2% of the total minimization time, and using the full code in ATM would make matters much worse.
After the preprocessing step, search algorithms (i.e., GA and NSGA-II) were then employed based on similarity values to find an optimal subset of test cases by removing redundant test cases. Using GA with combined similarity, ATM achieved a better trade-off in terms of effectiveness and efficiency compared to a SOTA test suite minimization approach, FAST-R [6]. However, it suffered from scalability issues as, for example, it took more than $10$ hours to minimize $4k$ test cases for the largest test suite. This motivated us to compare LTM to ATM, in terms of fault detection capability and, more importantly, minimization time. We aim to assess whether LTM can be more scalable than ATM while achieving comparable or even higher fault detection capability. We compared LTM to the two best configurations of ATM, using GA with combined similarity (ATM/Combined) and tree edit distance (ATM/TreeEditDistance). The former was the best configuration considering the trade-off between effectiveness and efficiency, whereas the latter achieved higher fault detection capability but was more time-consuming than the former in terms of total minimization time. We further evaluated the achieved saving in testing time for every minimized test suite for both LTM and ATM. We relied on the publicly available replication package of ATM [^6] https://doi.org/10.5281/zenodo.7455766 provided by its authors.